# Towards Harmonized Uncertainty Estimation for Large Language Models

### v1

@@ Abstract
Rui Li1, Jing Long1, Muge Qi1, Heming Xia2, Lei Sha3, Peiyi Wang1, Zhifang Sui1 1Peking University  2The Hong Kong Polytechnic University  3Beihang University o_l1ru1@stu.pku.edu.cn
@@ 1 Introduction
Harmonized uncertainty estimation is expected to encompass three key aspects: 1) Indication. The uncertainty score should clearly reflect the reliability of model responses, with higher scores signaling potential inaccuracies. This can be framed as a classification task, with “reliable” or “unreliable” as the classes. 2) Balance. Within classification framework, it’s critic to strike a balance between recall and precision, ensuring that challenging cases are appropriately flagged while minimizing the resources spent on false positives. 3) Calibration. The uncertainty score should align with human intuition and probabilistic expectations, to facilitate effective calibration. By striking a harmonized balance between these three aspects, uncertainty estimation provides an ideal measure of the model’s reliability, offering both usability and interpretability.
@@
In this paper, we introduce CUE[Uncaptioned image], a simple yet effective framework for adjusting uncertainty scores, which is orthogonal to existing uncertainty estimation methods. Specifically, we begin by curating dataset that are closely aligned with the target LLM’s performance within a particular domain of knowledge. This dataset is then utilized to train an auxiliary lightweight model, which serves as a Corrector to adjust the uncertainty scores. By integrating the Corrector trained on global alignment information with those uncertainty estimation methods that rely solely on the intrinsic logic and linguistic features of LLMs, we can significantly refine the uncertainty scores.
@@ 2.1 Uncertainty Estimation for LLMs
Logit-based methods are the most widely used and effective approaches in uncertainty estimation. Predictive Entropy (PE) (Malinin and Gales, 2020) defined uncertainty as the entropy of the output logits distribution, which is widely adopted and built upon in subsequent research. Follow that, Kuhn et al. (2023) introduced semantic entropy (SE) that estimates uncertainty by marginalizing over semantically-equivalent samples in NLG tasks. Duan et al. (2023) proposed Shifting Attention to Relevance (SAR), which focus on relevant information and assigns significance weights to tokens based on their contributions to the overall response. Yaldiz et al. (2024) introduced a Learnable Response Scoring Function (LARS), which utilizes supervised data to capture complex token-probability dependencies.
@@ 3.1 Limitation of Existing UE Methods
Some enhanced methods such as Length-normalized Predictive Entropy (LN-PE) (Malinin and Gales, 2020), SAR-t, SAR-s, SAR, and Semantic Entropy (SE) (Kuhn et al., 2023), make tailored adjustments to refine predictive entropy process, which show improvements over PE in terms of AUROC. However, no one is universally optimal for all target models and datasets. Moreover, as depicted in Figure 2 and Figure 5, the F1 scores of those methods are particularly low. This indicates that although those methods provided uncertainty scores with some potential to indicate the reliability of model response, they still fall short in strike a balance between precision and recall.
@@ 4 Method
In this section, we introduce CUE[Uncaptioned image], a correction framework featuring an intuitive approach to directly optimized for uncertainty estimation, where a Corrector is trained using a lightweight model to refine the uncertainty score. Through this method, we provide a more robust solution for uncertainty estimation. As shown in Figure 4, Our method comprises three main steps including dataset crafting, corrector training and uncertainty correcting.
@@ 4.2 Corrector Training
Employing the correction dataset $\mathcal{D^{*}}_{\text{cor}}$, we train a classifier to align with the performance of the target model. Specifically, the classifier integrates a fully connected layer following a lightweight encoder model, such as RoBERTa (Liu, 2019) and Deberta (He et al., 2021a), with the representation of the special token $[CLS]$ as its input, denote as $\mathbf{h}_{[CLS]}\in\mathbb{R}^{d}$ The output of the classifier is given by $\hat{y}_{i}=\sigma\left(\mathbf{W}\cdot\mathbf{h}_{[CLS]}+b\right),$ where $\sigma(z)$ is the sigmoid function, used to compute the likelihood $y_{i}$ that a data point belongs to label $c_{1}$. During training, we minimize the binary cross-entropy loss function $\mathcal{L}=-\sum{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$ across the correction dataset.
@@ 5.1.1 Models
We employed many lightweight encoder model as the base model to train the Corrector, including models from the RoBERTa series Liu (2019) and DeBERTa series He et al. (2021a, b).
@@ 5.2 Main Result
As illustrated in Table 1, the Corrector has brought in significant improvements, with an average AUROC score increase of 0.27 for TriviaQA and 0.09 for SciQA. Even when applied to challenging methods such as SE and SAR, the Corrector boosts AUROC scores by 0.01 to 0.03. Since AUROC reflects the UE method’s ability to assign higher scores to instances where the target model responds unreliably compared to those it responds to reliably, these improvements indicate that the deployment of the Corrector enhances the overall indicative capacity of the uncertainty scores, making it more effective for users in determining whether to trust the model’s responses.
Furthermore, as illustrated in Figure 5, the F1 score is also boosted by the Corrector, achieving an average increase of 38.97%. This notable improvement demonstrates the Corrector’s ability to help balance precision and recall, effectively mitigating the polarization tendency in the uncertainty scores observed in previous methods. We also verify the effectiveness of Corrector from a statistical perspective, see the LABEL:t-test for details.
@@
Although calibration is not the direct training objective of our Corrector, its application yields favorable calibration results. When employing the OPT-6.7B model as the target, we observed average ECE reductions of 0.34 on TriviaQA and 0.21 on SciQA. With the LLaMA-3-8B-Instruct model as the target, the reductions are 0.11 and 0.07, respectively—still considerable. To further illustrate the calibration performance facilitated by the Corrector, we provide calibration plots in Figure 6. We also conducted extensive experiments on the generalization performance of the Corrector. Due to space constraints, we put the detailed results and analysis in the AppendixLABEL:app:gene.
@@ 5.3 Ablation Study
We utilized various encoder models as base models to train the Corrector and assess the impact on correction performance. Specifically, we employed models from the RoBERTa series, including RoBERTa-base444huggingface.co/FacebookAI/roberta-base and RoBERTa-large555huggingface.co/FacebookAI/roberta-large, as well as models from the DeBERTa series, such as DeBERTa-base666huggingface.co/microsoft/deberta-base, DeBERTa-v3-large777huggingface.co/microsoft/deberta-v3-large, DeBERTa-v3-base888huggingface.co/microsoft/deberta-v3-base, DeBERTa-v3-small999huggingface.co/microsoft/deberta-v3-small, and DeBERTa-v3-xsmall101010huggingface.co/microsoft/deberta-v3-xsmall. These models represent different types, series, and sizes. As illustrated in Figure LABEL:fig:base_model, more advanced, later-generation, and larger models yield superior results.
@@
We compare correction scores from a lightweight classifier with those estimated using GPT-4o. We attempt not to rigorously assess the target model’s answers but to predict its reliability. Despite GPT-4’s strong performance in question answering, our results show it is less effective than the classifier in directly predicting reliability of target models when faced with questions. Additionally, as detailed in Section 3.2, combining uncertainty scores from different UE methods does not improve and may even degrade performance. This highlights the Corrector’s unique role as a complement to existing methods.
Other supervised methods for uncertainty estimation were also compared with the Corrector. The experimental results demonstrated that our supervised pipeline offered significant advantages in enhancing uncertainty estimation. Detailed results and analysis can be found in Appendix LABEL:app:comparative.

### v2

@@ Abstract
Rui Li1, Jing Long1, Muge Qi1, Heming Xia2, Lei Sha3, Peiyi Wang1, Zhifang Sui1thanks:   Corresponding author 1 Peking University 2 The Hong Kong Polytechnic University  3Beihang University o_l1ru1@stu.pku.edu.cn
@@ 1 Introduction
Harmonized uncertainty estimation is expected to encompass three key aspects: 1) Indication. The uncertainty score should clearly reflect the reliability of model responses, with higher scores signaling potential inaccuracies. This can be framed as a classification task, with “reliable” or “unreliable” as the classes. 2) Balance. Within classification framework, it’s critical to strike a balance between recall and precision, ensuring that challenging cases are appropriately flagged while minimizing the resources spent on false positives. 3) Calibration. The uncertainty score should align with human intuition and probabilistic expectations, to facilitate effective calibration. By striking a harmonized balance between these three aspects, uncertainty estimation provides an ideal measure of the model’s reliability, offering both usability and interpretability.
@@
In this paper, we introduce CUE[Uncaptioned image], a simple yet effective framework for adjusting uncertainty scores, which is orthogonal to existing uncertainty estimation methods. Specifically, we begin by curating dataset that is closely aligned with the target LLM’s performance within a particular domain of knowledge. This dataset is then utilized to train an auxiliary lightweight model, which serves as a Corrector to adjust the uncertainty scores. By integrating the Corrector trained on global alignment information with those uncertainty estimation methods that rely solely on the intrinsic logic and linguistic features of LLMs, we can significantly refine the uncertainty scores.
@@ 2.1 Uncertainty Estimation for LLMs
Logit-based methods are the most widely used and effective approaches in uncertainty estimation. Predictive Entropy (PE) (Malinin and Gales, 2020) defined uncertainty as the entropy of the output logits distribution, which is widely adopted and built upon in subsequent research. Following that, Kuhn et al. (2023) introduced semantic entropy (SE) that estimates uncertainty by marginalizing over semantically-equivalent samples in NLG tasks. Duan et al. (2023) proposed Shifting Attention to Relevance (SAR), which focuses on relevant information and assigns significance weights to tokens based on their contributions to the overall response. Yaldiz et al. (2024) introduced a Learnable Response Scoring Function (LARS), which utilizes supervised data to capture complex token-probability dependencies.
@@ 3.1 Limitation of Existing UE Methods
Some enhanced methods such as Length-normalized Predictive Entropy (LN-PE) (Malinin and Gales, 2020), SAR-t, SAR-s, SAR, and Semantic Entropy (SE) (Kuhn et al., 2023), make tailored adjustments to refine predictive entropy process, which show improvements over PE in terms of AUROC. However, no one is universally optimal for all target models and datasets. Moreover, as depicted in Figure 2 and Figure 5, the F1 scores of those methods are particularly low. This indicates that although those methods provided uncertainty scores with some potential to indicate the reliability of model response, they still fall short in striking a balance between precision and recall.
@@ 4 Method
In this section, we introduce CUE[Uncaptioned image], a correction framework featuring an intuitive approach to directly optimizing for uncertainty estimation, where a Corrector is trained using a lightweight model to refine the uncertainty score. Through this method, we provide a more robust solution for uncertainty estimation. As shown in Figure 4, Our method comprises three main steps including dataset crafting, corrector training and uncertainty correcting.
@@ 4.2 Corrector Training
Employing the correction dataset $\mathcal{D^{*}}_{\text{cor}}$, we train a classifier to align with the performance of the target model. Specifically, the classifier integrates a fully connected layer following a lightweight encoder model, such as RoBERTa (Liu, 2019) and Deberta (He et al., 2021a), with the representation of the special token $[CLS]$ as its input, denote as $\mathbf{h}_{[CLS]}\in\mathbb{R}^{d}$. The output of the classifier is given by $\hat{y}_{i}=\sigma\left(\mathbf{W}\cdot\mathbf{h}_{[CLS]}+b\right),$ where $\sigma(z)$ is the sigmoid function, used to compute the likelihood $y_{i}$ that a data point belongs to label $c_{1}$. During training, we minimize the binary cross-entropy loss function $\mathcal{L}=-\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$ across the correction dataset.
@@ 5.1.1 Models
We employed lightweight encoder models as the base model to train the Corrector, including models from the RoBERTa series Liu (2019) and DeBERTa series He et al. (2021a, b).
@@ 5.2 Main Result
As illustrated in Table 1, the Corrector has resulted in significant improvements, with an average AUROC score increase of 0.27 for TriviaQA and 0.09 for SciQA. Even when applied to challenging methods such as SE and SAR, the Corrector boosts AUROC scores by 0.01 to 0.03. Since AUROC reflects the UE method’s ability to assign higher scores to instances where the target model responds unreliably compared to those it responds to reliably, these improvements indicate that the deployment of the Corrector enhances the overall indicative capacity of the uncertainty scores, making it more effective for users in determining whether to trust the model’s responses.
Furthermore, as illustrated in Figure 5, the F1 score is also boosted by the Corrector, achieving an average increase of 38.97%. This notable improvement demonstrates the Corrector’s ability to help balance precision and recall, effectively mitigating the polarization tendency in the uncertainty scores observed in previous methods.
@@
Although calibration is not the direct training objective of our Corrector, its application yields favorable calibration results. When employing the OPT-6.7B model as the target, we observed average ECE reductions of 0.34 on TriviaQA and 0.21 on SciQA. With the LLaMA-3-8B-Instruct model as the target, the reductions are 0.11 and 0.07, respectively—still considerable. To further illustrate the calibration performance facilitated by the Corrector, we provide calibration plots in Figure 6.
We also conducted extensive experiments on the generalization performance of the Corrector. Due to space constraints, we put the detailed results and analysis in the Appendix LABEL:app:gene.
@@ 5.3 Ablation Study
We utilized various encoder models as base models to train the Corrector and assess the impact on correction performance. Specifically, we employed models from the RoBERTa series, including RoBERTa-base444huggingface.co/FacebookAI/roberta-base and RoBERTa-large555huggingface.co/FacebookAI/roberta-large, as well as models from the DeBERTa series, including DeBERTa-base666huggingface.co/microsoft/deberta-base, DeBERTa-v3-large777huggingface.co/microsoft/deberta-v3-large, DeBERTa-v3-base888huggingface.co/microsoft/deberta-v3-base, DeBERTa-v3-small999huggingface.co/microsoft/deberta-v3-small, and DeBERTa-v3-xsmall101010huggingface.co/microsoft/deberta-v3-xsmall. These models represent different types, series, and sizes. As illustrated in Figure 7, more advanced, later-generation, and larger models yield superior results.
Figure 7: The overall AUROC score gains achieved by Correctors trained on different base models across various UE methods on the SciQA dataset and Llama-3-8B-Instruct target model.
@@
We compare correction scores from a lightweight classifier with those estimated using GPT-4o. We attempt not to rigorously assess the target model’s answers but to predict its reliability. Despite GPT-4’s strong performance in question answering, our results show it is less effective than the classifier in directly predicting reliability of target models when faced with questions. Additionally, as detailed in Section 3.2, combining uncertainty scores from different UE methods does not improve and may even degrade performance. This highlights the Corrector’s unique role as a complement to existing UE methods.
We also conducted experiments to compare our Corrector with other supervised methods specifically designed for uncertainty estimation. The experimental results demonstrated that our supervised pipeline offered significant advantages in enhancing uncertainty estimation. Detailed results and analysis can be found in Appendix LABEL:app:comparative.
@@ 6 Conclusion
Acknowledgements
This paper is supported by NSFC project 62476009.